# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 16 Mar 2026 v4

@@ Abstract
Interestingly, the infrared response generated by a flat TDOS is structurally identical to anomalous dynamical susceptibilities that have appeared in several apparently unrelated contexts [5, 32]. In particular, the resulting spectrum corresponds in frequency space to a logarithmic real part and a nearly constant absorptive component of the susceptibility. Similar infrared structures were introduced phenomenologically in the marginal Fermi liquid theory of strange metals and also arise in solvable strongly correlated spin models such as the Sachdev–Ye spin liquid. In the present framework, however, this behavior follows directly from the accumulation of slow relaxation modes and therefore emerges naturally as a consequence of a finite infrared relaxation-rate density of states.
The lower panel of Fig. 4 illustrates the direct consequence of this spectral organization for superconductivity. Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic Cooper kernel $\Pi(T)\sim\ln(\Omega/T)$ and hence to the exponentially small transition scale characteristic of conventional marginal pairing theories. In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$. The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
@@
A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order is governed by condensation in a Cooper sector embedded within an extended slow-mode manifold rather than by binding through a sharp bosonic excitation.
@@
A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena. The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
@@
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a generic mathematical feature of dynamical systems with extended temporal correlations rather than a model-specific construction. Related derivations are presented in Appendix F.
This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics. While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem continuously exchanging coherence and information with the remaining degrees of freedom. From the perspective of such a subsystem, the relevant generator of dynamics is not the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes. The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues rather than solely by the spectrum of excitation energies. In this sense the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.
From this perspective Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
@@
An important conceptual aspect of the present framework concerns the role of temporal scaling in dynamical criticality. In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges indirectly through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$. The Hertz–Millis theory of quantum critical metals follows a similar philosophy, in which spatial critical fluctuations generate an effective dynamical scaling once fermionic degrees of freedom are integrated out.
@@
In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves. When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales. The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions, with logarithmic or stronger singular structures depending on the observable sector. Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality. This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena [33, 34].
@@
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. A decisive control parameter for high transition temperatures may be the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this viewpoint certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat-TDOS regime.
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [35]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. In this perspective high-temperature superconductivity is most naturally favored in regimes characterized by maximal infrared memory rather than simply by extreme coupling strength.
@@
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting marginal logarithmic instabilities into algebraic superconducting transitions governed by infrared spectral weight.
@@
In this perspective, high–temperature superconductivity can be viewed as a generic consequence of favorable infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-K_{R}(t-t^{\prime})\right]$ (S67)
@@
where $\tilde{\phi}$ denotes the response field, $K_{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
@@
$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&K_{A}(\omega)\\ K_{R}(\omega)&iK_{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S69)
where $K_{R,A}$ are retarded/advanced kernels and $K_{K}$ is the Keldysh (noise) component. The retarded kernel encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
@@
and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $K_{K}$. The corresponding MSRJD action then takes the form (S67), with the correspondence
$K_{R}(\omega)\;\leftrightarrow\;K_{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}K_{K}(\omega).$ (S71)
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
Appendix D Appendix D: Derivation of the Cooper-channel polarization kernel from a relaxational mode continuum
In this Appendix we present an explicit derivation of the polarization kernel used in the Cooper channel within the relaxation-spectrum framework. The Cooper-channel object relevant for pairing enhancement is the fluctuation (or correlator) spectrum of the collective pair field.
The construction below should be understood, in the same spirit as Appendix B, as an effective infrared description after microscopic electronic degrees of freedom have already been integrated out. At that stage the relevant dynamical variables are collective pair modes, and the polarization kernel is obtained from their stochastic two-point correlations rather than from an explicit microscopic fermionic loop.
(I) Single relaxational pair mode: Consider a single complex collective pair mode $\Delta_{\lambda}(t)$ with relaxation rate $\lambda>0$ obeying the linear Langevin equation

### 30 Mar 2026 v5

@@ Abstract
Interestingly, the infrared response generated by a flat TDOS is closely related to anomalous dynamical susceptibilities that have appeared in several apparently unrelated contexts [5, 34]. In particular, the resulting spectrum corresponds to a logarithmic real part and a weakly frequency-dependent absorptive component of the susceptibility. Similar infrared structures were introduced phenomenologically in the marginal Fermi liquid theory of strange metals and also arise in solvable strongly correlated spin models such as the Sachdev–Ye spin liquid. In the present framework, however, this behavior follows directly from the accumulation of slow relaxation modes and therefore emerges naturally as a consequence of a finite infrared relaxation-rate density of states.
The lower panel of Fig. 4 illustrates the consequence of this spectral organization for superconductivity. Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, the transition scale is controlled by the infrared structure of the pairing kernel. In conventional Ohmic or marginal pairing scenarios, the Cooper kernel acquires a logarithmic enhancement $\Pi(T)\sim\ln(\Omega/T)$, leading to the exponentially small transition scale characteristic of BCS-like theories. In the memory-dominated regime with a flat TDOS, the retarded pairing kernel similarly exhibits logarithmic infrared enhancement, $\Pi(T)\sim\rho_{0}\ln(\Lambda/T)$. While the transition scale retains a BCS-like exponential form, the effective parameters are now controlled by the infrared spectral weight of slow collective modes. Only when the TDOS becomes infrared-singular ($\alpha<0$) does the pairing kernel acquire a power-law enhancement, $\Pi(T)\sim T^{\alpha}$, leading to algebraic scaling of the transition temperature.
@@
A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. In the marginal regime, the exponential suppression characteristic of weak-coupling BCS theory is not removed but can be substantially reduced due to enhanced infrared spectral weight. Only when the TDOS becomes infrared-singular does the pairing scale exhibit algebraic enhancement. In this sense, superconducting order is governed by condensation in a Cooper sector embedded within an extended slow-mode manifold rather than by binding through a sharp bosonic excitation.
@@
A single infrared parameter, $\rho_{0}$, plays a central role in both thermodynamic and dynamical phenomena. The superconducting dome can be interpreted as reflecting truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
@@
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel formulations admit representations in terms of decay modes characterized by relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations naturally lead to continuous relaxation spectra. A dense TDOS can therefore be understood as a generic feature of dynamical systems with extended temporal correlations rather than a model-specific construction.
This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics. While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem exchanging coherence and information with the remaining degrees of freedom. From this perspective, the relevant generator of dynamics is not the Hamiltonian itself but an effective Liouvillian governing relaxation and memory processes. The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues in the effective description.
In this sense, the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.
@@
An important conceptual aspect of the present framework concerns the role of temporal scaling in dynamical criticality. In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$. The Hertz–Millis theory of quantum critical metals follows a similar philosophy, in which spatial critical fluctuations generate an effective dynamical scaling once fermionic degrees of freedom are integrated out.
@@
In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves. When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales. The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
Scale-free temporal dynamics can thus be viewed as emerging directly from the structure of the relaxation spectrum, rather than being solely inherited from spatial criticality. This perspective is closely related to scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural framework for describing dynamical critical phenomena [35, 36].
@@
Memory–dominated dynamics and prospects for high-temperature superconductivity. Within the present framework, superconducting transition temperatures are controlled not only by the microscopic pairing interaction but also by the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
In the marginal regime corresponding to a flat TDOS ($\alpha=0$), the retarded pairing kernel exhibits logarithmic infrared enhancement, leading to a BCS-like transition scale, $T_{c}\sim\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right)$, where $\Lambda$ is an electronic cutoff scale and $\rho_{0}$ is the infrared spectral weight of slow collective modes. Because $\Lambda$ is set by electronic energy scales, this form naturally allows transition temperatures in the range of $10^{2}$ K for moderate coupling, providing a consistent framework for understanding high-$T_{c}$ superconductivity without invoking finely tuned bosonic mediators.
Further enhancement of the transition scale may occur if the relaxation spectrum develops an infrared-singular structure ($\alpha<0$), corresponding to an accumulation of slow modes toward vanishing relaxation rates. In this regime, the pairing kernel acquires a power-law infrared enhancement, and the transition temperature follows an algebraic scaling with the effective interaction strength.
These considerations suggest a complementary perspective on superconductivity design: beyond increasing microscopic pairing interactions, engineering the infrared organization of collective dynamics may provide an effective route toward higher transition temperatures [37]. In this sense, dynamical phases supporting enhanced slow-mode spectral weight represent promising candidates for realizing superconductivity at elevated temperatures.
@@
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically enhanced, leading to logarithmic infrared amplification of the retarded pairing kernel. While the resulting transition scale retains a BCS-like exponential form, it is controlled by the infrared spectral weight of slow collective modes. Algebraic enhancement of the transition temperature arises only when the TDOS becomes infrared-singular ($\alpha<0$), corresponding to an accumulation of slow modes toward vanishing relaxation rates.
@@
In this perspective, high–temperature superconductivity can be understood as arising from favorable infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-L^{R}(t-t^{\prime})\right]$ (S67)
@@
where $\tilde{\phi}$ denotes the response field, $L^{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
@@
$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&L^{A}(\omega)\\ L^{R}(\omega)&iL^{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S69)
where $L^{R,A}$ are retarded/advanced propagators and $L^{K}$ is the Keldysh (fluctuation) component. The retarded component encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
@@
and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $L^{K}$. The corresponding MSRJD action then takes the form (S67), with the correspondence
$L^{R}(\omega)\;\leftrightarrow\;L^{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}L^{K}(\omega).$ (S71)
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
Appendix D Appendix D: Retarded Keldysh derivation of the superconducting transition scale from the collective Cooper response
In this Appendix we reformulate the Cooper-channel instability directly in the real-time Keldysh language. The central point is that the superconducting transition temperature must be determined from the retarded linear-response kernel of the collective pair field, rather than from the Keldysh fluctuation correlator itself. The fluctuation propagator will be analyzed subsequently, but the transition criterion is fixed first by the retarded ladder pole.
We therefore proceed in the following order. First, we introduce the Keldysh quadratic action for the collective pair field. Second, we derive the retarded propagator of a single relaxational pair mode. Third, we integrate over a continuum of relaxational modes with a TDOS characterized by a general infrared exponent $\alpha$. Fourth, we resum the retarded Cooper ladder and derive the Thouless criterion for $T_{c}$. Finally, we analyze explicitly the cases $\alpha=0$ and $\alpha=-1/2$.
(I) Collective pair field and quadratic Keldysh action: We begin from the collective Cooper pair field
$\Delta(x,t)=c_{\uparrow}(x,t)c_{\downarrow}(x,t),$ (S101)
and its Keldysh classical/quantum components
$\Delta_{cl}(x,t)=\frac{1}{2}\left(\Delta_{+}(x,t)+\Delta_{-}(x,t)\right),$ (S102)
$\Delta_{q}(x,t)=\Delta_{+}(x,t)-\Delta_{-}(x,t),$
where ++ and - denote the forward and backward Keldysh contours.
At quadratic level, the real-time action of the collective pair sector has the generic form
$S_{2}=$ (S103)
$\int_{\omega}\begin{pmatrix}\Delta_{cl}^{*}(\omega)&\Delta_{q}^{*}(\omega)\end{pmatrix}\begin{pmatrix}0&\left[L_{0}^{A}(\omega)\right]^{-1}\\ \left[L_{0}^{R}(\omega)\right]^{-1}&\Sigma^{K}(\omega)\end{pmatrix}\begin{pmatrix}\Delta_{cl}(\omega)\\ \Delta_{q}(\omega)\end{pmatrix}$,
where $\int_{\omega}\equiv\int d\omega/(2\pi)$. The corresponding propagator matrix is
$\hat{L}(\omega)=\begin{pmatrix}L^{K}(\omega)&L^{R}(\omega)\\ L^{A}(\omega)&0\end{pmatrix}$, (S104)
with
$L^{R}(\omega)=\frac{1}{\left[L_{0}^{R}(\omega)\right]^{-1}},\quad L^{A}(\omega)=\frac{1}{\left[L_{0}^{A}(\omega)\right]^{-1}}.$ (S105)
The superconducting instability is determined by the pole of the retarded pair susceptibility. Therefore, throughout the first part of this Appendix, the relevant object is the retarded pair response kernel.
(II) Retarded propagator of a single relaxational pair mode: Consider first a single collective pair mode $\Delta_{\lambda}(t)$ with relaxation rate $\lambda>0$. Its linear relaxational equation is